

********************************************************************
UARS WEEKLY STATUS REPORT - 26 MAR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems. The data from the new working channel show that it
is working as expected.
Until the recent changeout of the elements of the working (daily) optical
channel, the V19r3 MgII index was generated only days after data was
taken. This has been temporarily suspended until the effects of the
change are more thoroughly analyzed. We expect to resume near-current
day generation of the MgII index within a month or so.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
PEM is active under PMON control. Energy processor #6 for the
HEPS-1 unit has been found to not activate correctly. Data affected
are from the upper energy EE channel for Head-2 of HEPS-1 (that is the
+15 deg electrons above 300 keV). Tests are being conducted on this
energy processor and other energy processors to determine conflicts.
Measurements from this sensor will be activated as soon as possible.
The PEM team and FOT are convinced that the entanglement of energy
processor #6 is related to OBC control and not a failure of that
energy processor. Early attempts at adjusting OBC command timing
have not been successful. The FOT and the PEM team is attempting to
refine their diagnostics and confine the problem.
While under PMON control, the PEM instrument computes energy deposition
from the particle spectrometers.
***********
HALOE
***********
HALOE was turned off at the end-of-day March 24, 1999 for the upcoming yaw.
HALOE had been successfully operating up to that point. HALOE is
scheduled to be turned on near the end-of-day March 31, 1999. There has
been no detectable change in the performance of the science channels.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII reprocessing is well under way with Version V5_11, running at a
rate of approximately 9 days of data per day.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
6 days CLA (7.0) ENG (L1-L3) (6R-6C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-328C-1F=431L)
126 days SUS (7.0) PROD (677R-665C=12L)
92 days WIN (5.11) PROD (LOG) (2014R-223C=1791L)
0 days WIN (5.11) PROD (L1-L3) (2014R-13C=2001L)
The CDHF sent out UCSS/UOAS/NRT Release Notes for this recently installed
Y2K compliant software.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
SOLSTICE Wall Clock time increase.
Efforts to run the debug version of the SOLSTICE job step 2 program on
the "Alpha" had been halted as a result of the "Insufficient virtual
memory" error. The VAX version, completely recompiled and relinked,
now failed for the same reason. More debug was used to determine why
the calculation for the number of bytes required for a DBI_FULL_REQUEST
message resulted in a value of 2GB of memory! Additional debug cleared up
the problem. This may indicate a compilation switch problem in the code.
The VAX (debug) version is now able to run--although not to completion.
***********
FOT
***********
This report covers 13 March 1999 (Orbit 40990, GMT Day 072) through 19 March
1999.
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 075/2110 41048 AXIS Off
WINDII 075/2110 41048 Off
The beta angle ranged between the angles of 51.3 degrees to 33.4 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 28 March 1999 (DOY 087).
A +0.5 to 0.0 Degree Roll Reset Maneuver was performed on 16 March 1999
(DOY 075) in Orbit 41021 at 075/23:38:06. The WINDII instrument had
finished its observation of Polar Mesospheric Clouds and UARS returned to
level flight.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 12 March 1999 and 19 March 1999 are provided in
Appendix D.
The clock error ranged from -11.8 to 10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 02
% PEM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496948
Transponder B: 99/055 - 2287.498731
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/077 show the observatory data loss to be
22 hours, 46 minutes, 35.028 seconds (an increase of 2.272 seconds since
last report period). This is a 0.0347 percent data loss which equals a
99.9653 percent data capture for the mission.
The increase of 2.272 seconds occurred over two (2) orbits:
7 1.152 seconds on Day 075 due to no overlap, and
7 1.120 seconds on Day 077 due to data hit (IP).
One additional Y2K "test" 8-week planning aid was generated by FDF this
week. Unfortunately, this file (the S/C Orbital Events file) contained
some sort of formatting problem which did not allow it to pass through the
FORMATS system. The FOT is still awaiting the S/C Orbital Events file, as
well as the Ground Track planning aid file.
The UARS Test and Training Simulator (UTTS) has been repaired and was
declared operational on 18 March 1999. Tests will be performed to verify
correct operation.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. PEM EP6 is still not coming up nominally under
PMON control. FOT continued diagnostic procedures under direction from R.
Frahm, SWRI. PDB-10 was disabled on Wednesday, 17 March. EP-6 remained
enabled when the three PDB-10 commands were inserted. A command sheet to
restore PDB-10 activation was created on Friday afternoon, 19 March to be
loaded shortly after the change of day. After SWRI analyzes the most
recent test data, we anticipate further direction on troubleshooting the
EP6 activation.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
MEETINGS
FOT supported an IMOC meeting on 17 March 1999. The organizational chart
for the IMOC integration and development teams were presented as well as
the deadlines for various reports and presentations that must be written or
presented.
FOT supported an UTTS meeting on 19 March 1999. The UTTS problem, fix and
current status were discussed. The persons responsible for the different
parts of the UTTS were identified. New reporting and accountability
procedures were discussed to help prevent communication problems when
troubleshooting future problems.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 28
March 1999 (DOY 087).
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 March 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41027 074/1025 99-103 OBC Clk Adj fr 20.6 to 0.00 msec/day
41048 075/2110 99-104 PEM AXIS Power Off
41048 075/2110 99-105 WINDII Power OFF
41052 076/0236 99-106 HRDI Power ON HIGH
41053 076/0423 99-107 HRDI Power ON LOW
41059 076/1432 99-108 OBC Clk Adj fr 0.00 to 20.6 msec/day
41060 076/1622 99-109 PEM PEM Test
41086 078/0904 99-110 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 March 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 March 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41027 074/10:30 (15 Mar 99) from 20.6 to 00.0 msec/day
41059 076/14:36 (17 Mar 99) from 00.0 to 20.6 msec/day
41086 078/09:06 (19 Mar 99) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 March 1999
12 Mar 1999 - GMT Day 071 Beta = 51.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -22.4/-44.8 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-5.6 +5.47/+4.51 +26.8/-15.6 25.0 1.022 20.8
3 +56.0/-5.6 +4.51/+3.23 +26.4/-16.0 25.0 1.022 20.6
19 Mar 1999 - GMT Day 078 Beta = 33.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -22.4/-44.8 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +56.0/ 0.0 +5.15/+3.87 +29.6/-14.4 25.9 1.025 20.6
3 +44.8/-22.4 +3.87/+2.60 +28.8/-14.4 25.9 1.019 20.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 March 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-012 FOT 40552 042/1739 20681 Data Loss (38:18)
99-011 FOT 41068 077/0351 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-012 FOT 40552 042/1739 20681 Data Loss (38:18)
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
99-012 FOT 40552 042/1739 20681 Data Loss (38:18)
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-011 FOT see list Mar 99 14170 Generic Late Acq
SEVNTFW99/12